DIFFERENTIAL GEOMETRY/PDE SEMINAR

Wednesday, January 26, 2005 Padelford C-36 3:50 pm

Radiation fields, scattering and inverse scattering on asymptotically hyperbolic manifolds

Antonio Sa Barreto (Purdue University)

We will discuss the radiation fields, and how they effectively substitute the Dirichlet to Neumann map in this setting. We will then show how the boundary control method can be adapted to show that the scattering matrix at all energies determines the manifold and the metric, modulo certain invariants.

If time permits, we will discuss the difference between this and (the more difficult) asymptotically Euclidean case.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.